OpenStudy (anonymous):

Find the derivatives using the table (I will post the table) a. Dx(f[g(x)]) at x=1 b. same at x=2

OpenStudy (anonymous):

x 1 2 3 4 f(x) 2 4 1 3 f'(x) -6 -7 -8 -9 g(x) 2 3 4 1 g'(x) 2/7 3/7 4/7 5/7

OpenStudy (anonymous):

∫ f ′(x) f (x) dx = ln ( f (x)) + C

OpenStudy (bahrom7893):

(f[g(x)])' = f '(g(x)) * g' (x), at x = 1 f ' (g(1)) * g'(1) = f ' (2) * (2/7) = -7 * 2/7 = -2

OpenStudy (bahrom7893):

that was the answer to the first one.

OpenStudy (anonymous):

thats for A

OpenStudy (anonymous):

awesome, thats what I got for the first one- not doing integrals yet so i cant use madiners

OpenStudy (bahrom7893):

for the second one: f ' (g(2)) * g'(2) = f ' (3) * 3/7 = -8*3/7 = -24/7

OpenStudy (anonymous):

for B x ln(x)d

OpenStudy (anonymous):

lol thanks bahrom...no idea what is going on in your answer madiner but thanks for the input

OpenStudy (anonymous):

thnk bahrom i didnt know that he didnt learn intergrels